Find two numbers whose sum is 58 and whose difference is 16.
Answer by pwac(253)

First number=x
Second number=x+16
So x+x+16=58
2x+16=58 subtract 16 from both sides
2x=42 divide both sides by 2
x=21
x+16=37
two numbers are:21 and 37
Pete